# Simulating Classroom Education with LLM-Empowered Agents

### v1

@@ 4.4 User Experience
In this section, we report the results from the student experience with SimClass. As shown in Figure 3, several key findings are observed: (1) Importance of Interactions. Interactions during class are crucial for users. Without interaction, user experience significantly declines across all three metrics. (2) Enhancement by Classmate Agents. Classmate agents enhance user experience in terms of Cognitive Presence and Social Presence. This enhancement may be attributed to the classmate agents’ active engagement in asking questions to the teacher, which aids the user’s understanding of concepts and increases the sense of Social Presence in the classroom. (3) Satisfying Teaching Presence. All systems demonstrate good Teaching Presence, maintaining a focused and coherent class. This metric largely depends on the quality of the teaching scripts used, though we observe that interaction and Student Roles slightly improve the user experience. (4) Better Experience in HSU with students. The HSU course achieves a better user experience with the full setting. HSU focuses on college interpersonal relationships and learning methods, where peer learning plays a more significant role. This suggests that a multi-agent design is particularly crucial for certain types of classes.
Conclusion of user experiments. According to Figure 2 and Figure 3, the classroom demonstrates the effectiveness in SimClass in terms of both interactions and user presence.
@@
| [Teacher]: Let’s talk about Alan Turing. … The concept of a thinking machine he proposed essentially foresaw many characteristics of modern intelligent systems, including the ability to learn, adapt, and handle complex problems. |
| [Classmate]: I wonder how Turing would react if he were alive today and saw the voice assistants on our smartphones. But he might also remind us that true intelligence is not merely about mimicking human behavior; it must involve its own “thinking” and “understanding” … |
@@
| [Teacher]: Of course, let me explain it to you. … When a neural network forward propagates data and predicts outputs during training, … we need to adjust the weights in the neural network. This adjustment process is carried out through backpropagation. … |
| [Classmate]: Thank you for sharing! I would like to share my notes. When we train the model, … |
@@
| [Classmate]: Hey, don’t think that way. Remember those smart house assistants? They didn’t know much at first, but how did they gradually learn? Through continuous learning and adaptation! We are the same. … |
@@
| [Classmate]: Classmate, hearing you say that suddenly made me think of an AI application. Maybe in the future, a chef robot could perfectly cook a steak based on your taste and health needs! Back to our topic, regarding AGI, … |
| [Teacher]: Yes, maybe we can imagine a future smart kitchen equipped with an AI system that customizes the cooking method based on your preferences. … |
@@ 4.5 Agent Behaviors
Based on our classification of various types of classroom interactions in Section 3, we present some emergent group behaviors observed during the classroom experiments in SimClass .
Teaching and Initiation. When the user learns from teaching, classmates engage and share their inspiring ideas, which deepens the depth of the topic and enriches the discussion. The diversity of agents, each approaching from different perspectives, introduces a wider range of possibilities for classroom teaching content.
In-depth Discussion. If the current explanation is not clear enough for the users, they can ask questions at any time to initiate a discussion with the teacher and classmates until clarity is achieved. This highlights the advantage of SimClass as an interactive classroom compared to one-to-many education methods like pre-recorded videos.
Emotional Companionship. Beyond knowledge dissemination, maintaining a positive learning atmosphere is crucial in classroom scenarios. When a user expresses negative learning intent, the classmate agent intervenes after the assistant, utilizing class content in the history and providing vivid emotional support as a non-teacher role.
Classroom Management. Similarly, when a user tries to interrupt the system, the classmate agent subtly redirects the class while following the user’s words. These classmates enhance classroom discipline more effectively than the teacher alone, demonstrating emergent group behaviors.
Conclusion of case study. Based on the cases above, We can observe diverse interactions between different class roles within the classroom, as well as the effectiveness of the manager agent, who is spontaneously capable of designating appropriate speakers to elicit emergent group behaviors of Class Roles seamlessly, which significantly enlivens the class and enhances the user’s experience.
@@ 5 Conclusion
We introduce SimClass, a novel framework of multi-agent classroom using LLMs to answer several fundamental research questions in the era of LLM-driven education. Based on several theoretical methods, our experiments span two courses with real users, demonstrating interaction patterns similar to those in real classrooms and effective learning experience in SimClass. We observe emergent collaborative behaviors among LLM agents during the teaching process. Future work could incorporate more agents and explore more courses to analyze more diverse classroom behaviors. We hope our efforts can advance the explorations of LLM-empowered systems for AI-driven education researchers, practitioners, and pedagogues.
@@ 6 Limitations
Despite our analytical efforts covering major theories in education, our work still has the following limitations: Firstly, we apply GPT-4 as our backbone model to perform our experiments. a more comprehensive understanding of our framework requires a broader range of diverse experiments. Secondly, we conduct experiments on a limited number of agents, while a more diverse set of agent characters could capture a wider array of behaviors in the classroom. Thirdly, we apply a limited quantity of functions in our system, while more various functions in teaching scenarios could further enhance the performance of the system.
@@ 7 Ethical Considerations
Our investigation involves the development of a simulated classroom environment populated by artificial intelligent models acting as classmates and teachers. All user data obtained throughout these interactions will be anonymized to ensure privacy and confidentiality. Informed consent is obtained from participants, who are thoroughly briefed on the nature of simulation, the AI generated content, and the data collection process. Participants receive appropriate compensation for their involvement.
In educational systems involving large language models, there is a potential for generating hallucinations and incorrect information. Therefore, applying these systems to real-world scenarios requires careful consideration and thorough evaluation before serving real users.
Appendix A Survey and Quiz
@@
In this appendix section, we present detailed designs of the surveys and quizzes in our experiments.
@@
Table 6 illustrates how the surveys were structured to evaluate three crucial dimensions of the learning experience: cognitive presence, teaching presence, and social presence. Each dimension includes rating guidelines to ensure consistent and reliable feedback from diverse users.
@@
The quizzes, administered after participants engaged with the simulated classrooms, were designed to measure the depth of their learning. The TAGI quiz evaluates understanding of artificial intelligence concepts, focusing on symbolic intelligence, pre-trained language models, and emergent phenomena (Table 8). Similarly, the HSU quiz assesses broader educational strategies and personal development, covering topics such as internal motivation, academic stress management, and time management (Table 7). All questions were meticulously crafted and verified by subject matter experts to align closely with course materials.
@@
Both quizzes feature multiple-choice questions, some with multiple correct answers, to comprehensively test whether the participants are actively engaged in the experiment.
Appendix B Examination of GPT-4 Labeling
To validate the GPT-4 labeling in our experiment, we sampled 100 data points labeled by GPT-4 and had an expert familiar with FIAS label them for comparison. The results showed that GPT-4’s labels matched the human expert’s labels with an accuracy of . We believe this demonstrates that GPT-4 can serve as a reliable and balanced alternative to crowd-sourced human labelers in our experiments. Additionally, we examined the eight instances where GPT-4’s labels differed from the human expert’s labels. These cases were also found to be uncertain during human labeling, suggesting that GPT-4 not only avoids individual human biases but also achieves a high level of precision comparable to human-labeled results.
Appendix C FIAS Matrices for Ablation Systems
In addition to the default setting of SimClass, we also provide the sum of the matrices based on Flanders Interaction Analysis System for our ablation settings (w/o classmate agents and w/o interactions), as demonstrated in Figure 4 and Figure 5. In comparison with Figure 2, different types of classes demonstrate different interaction patterns. Generally, the fewer types of interactions there are, the less diverse the classroom will be. This indicates the significance of adding more kinds of agents for interactions, especially classmate agents to simulate a vivid classroom.
| A. Proprietary Intelligence |
| B. Symbolic Intelligence |
| C. General Intelligence |
| D. Neural Network Intelligence |
| A. Masked Language Model |
| B. Next Sentence Prediction |
| C. Possibility Memorization |
| D. Next Token Prediction |
| A. Self-supervised Pre-training |
| B. Supervised Fine-tuning |
| C. Reinforcement Learning from Human Feedback |
| D. Instruction Tuning |
| A. In-context Learning |
| B. Chain-of-Thought |
| C. Sentiment Analysis |
| D. Instruction Following |

### v2

@@ 4.3.2 Community of Inquiry Analysis
In this section, we report the survey results in the ablation systems. As shown in Figure 4, several key findings are observed: (1) Interactions during class are crucial for users. Without interaction, user experience significantly declines across all three metrics. (2) Classmate agents enhance user experience in terms of Cognitive Presence and Social Presence, by actively engaging with the teacher, helping users understand concepts, and active discussions. (3) All systems maintain good Teaching Presence with focused, coherent classes, primarily influenced by the quality of the teaching scripts. Interaction and classmates further enhance the user experience. (4) The full multi-agent setup provides a better experience in HSU, which emphasizes college interpersonal relationships and learning methods, highlighting the importance of peer learning and multi-agent design for certain course types.
@@
| [Teacher]: Let’s talk about Alan Turing. … The concept of a thinking machine essentially foresaw many characteristics of modern intelligent systems, including the ability to learn, adapt, and handle complex problems. |
| [Deep Thinker]: I wonder how Turing would react if he were alive today and saw the voice assistants on our smartphones. But he might also remind us that true intelligence is not merely about mimicking human behavior; it must involve its own thinking and understanding … |
@@
| [Teacher]: Of course, let me explain it to you. When a neural network forward … we need to adjust the weights in the neural network. This adjustment process is carried out through backpropagation. … |
| [Note Taker]: Thank you for sharing! I would like to share my notes. When we train the model, … |
@@
| [Class Clown]: Hey, don’t think that way. Remember those smart house assistants? They didn’t know much at first, but how did they gradually learn? Through continuous learning and adaptation! We are the same. … |
@@
| [Inquisitive Mind]: Classmate, what you said made me think of an AI application. Maybe in the future, a chef robot could perfectly cook a steak based on your taste and health needs! Back to our topic, regarding AGI, … |
| [Teacher]: Yes, maybe we can imagine a future smart kitchen equipped with an AI system that … |
@@ 4.4 Agent Behaviors
Based on our classification of classroom interactions in Section 3, we present key group behaviors observed during the experiments in Table 5:
Teaching and Initiation. When the teacher teaches, classmates actively engage by sharing ideas, enriching discussions and deepening the topic. The agents’ diverse perspectives broaden the scope of the teaching content.
In-depth Discussion. When users seek clarity, they can ask questions, initiating interactive discussions with teacher and classmates. This highlights the strength of SimClass over one-to-many education methods like pre-recorded videos.
Emotional Companionship. Beyond knowledge dissemination, maintaining a positive learning atmosphere is crucial in classrooms. When a user expresses negative learning intent, the classmate agent intervenes after the assistant, utilizing class content in the history and providing vivid emotional support as a non-teacher role.
Classroom Management. When a user attempts to disrupt the class, the classmate agent gently redirects the session, acknowledges the user’s input, and hands control back to the teacher. This collaborative approach to maintaining order is more effective than teacher efforts alone.
These cases illustrate diverse interactions among class roles and the effectiveness of the Session Controller, which seamlessly designates appropriate speakers to encourage group behaviors, enhancing engagement and enriching the user experience.
@@ 5 Conclusion
We introduce SimClass, a novel multi-agent classroom framework leveraging LLMs for teaching. Our experiments across two courses with real users demonstrate its effectiveness in simulating dynamic teaching environments, where agents collaborate to enhance user experience. Increased interaction with the system results in better learning outcomes, and the multi-agent setup encourages students to engage more. We hope our efforts can advance the explorations of LLM-empowered education systems for researchers, practitioners, and pedagogues.
@@ 6 Limitations
System: The system we The system we are currently developing has several limitations, since it represents an initial exploration in this field. First, the manager agent and class roles are implemented using LLMs, which introduces response delays, particularly in scenarios where multiple agents need to participate. This can affect the overall user experience. Future work could focus on replacing our current design with higher-performance models to address this issue. Second, our framework requires designed slide-script pairs by teachers. Future efforts could aim at automating this process. Lastly, our system incorporates a limited set of teaching functions, which restricts its performance. Future developments could introduce more diverse forms of classroom interactions, drawn from educational settings, and integrate additional technologies to enhance the experience. For instance, retrieval-augmented generation (RAG) could be employed to improve knowledge accuracy, while question generation and knowledge tracing could be used to personalize the agents’ responses further.
Experiments: Due to cost and time constrains, our experiments were conducted with a limited number of courses, models, quizzes, and users. A more comprehensive evaluation of the framework would require a broader and more diverse set of experiments. As such, our findings may be constrained by the specific course types we used, the model available to us, and users we recruited. We hope that our work serves as a contribution to the broader discussion on the use of LLMs in simulating classroom roles for education. Further experiments should explore a wider variety of courses (across different subjects and levels of difficulty), a more diverse set of agents (with different personas, teaching strategies, and group sizes), additional quizzes for a more thorough analysis of learning outcomes, a larger group of users from different backgrounds, and a broader range of LLMs.
Participants: Our current experiments are conducted within the scope of general courses at the university level, focusing on college students. For a given course level, the abilities and proficiency of students tend to be similar, which introduces biases due to the homogeneity of the participant group. In future work, we aim to extend our system to benefit a broader range of users, with a particular focus on marginalized groups and individuals with learning disabilities, promoting greater educational equity.
@@ 7 Ethical Considerations
Our investigation involves the development of a simulated classroom environment populated by artificial intelligent models acting as classmates and teachers. All user data obtained throughout these interactions will be anonymized to ensure privacy and confidentiality. Informed consent is obtained from participants, who are thoroughly briefed on the nature of simulation, the AI generated content, and the data collection process. Participants receive appropriate compensation for their involvement. In educational systems involving LLMs, there is a potential for generating hallucinations and incorrect information. Therefore, applying these systems to real-world scenarios requires careful consideration and thorough evaluation before serving real users.
On the other hand, multi-agent teaching systems may lead to a different student perception of the role of teacher, comparing traditional classroom teachers. In the past, teachers were real individuals who adhered to social norms, whereas AI teachers focus more on knowledge delivery. Therefore, these systems may also have a bias in the development of students’ abilities.
While the agents in the classroom can enhance the learning experience, they cannot replace the role of human teachers in fostering students’ comprehensive skills, nor the role of real students in improving social skills, sense of group identity, and fostering self-esteem. Therefore, the use of these systems requires more diverse and interdisciplinary research, particularly with the guidance and input from fields like psychology and education.
Appendix A Experiment Details
For the reproducibility, we provide details of our online system and experiments.
Model Parameters. Regarding the model APIs we used, the online system utilized GLM-4, with the model named glm-4. For the ablation systems, we employed gpt-4-vision-preview, and for the FIAS classification, we used gpt-4-turbo. All models were run with the default temperature settings.
FIAS. For FIAS, due to the differences between our online system and traditional classroom settings, especially the difficulty in defining and measuring the "silence" metric, we excluded the category of silence. Consequently, for a fair comparison, in Table 2, we also removed the silence metric from the novice and expert teacher classrooms in the study by Zhang et al. (2023).
Examination of GPT-4 Labeling. To validate the GPT-4 labeling in our experiment, we sampled 100 data points labeled by GPT-4 and had an expert familiar with FIAS label them for comparison. The results showed that GPT-4’s labels matched the human expert’s labels with an accuracy of . We believe this demonstrates that GPT-4 can serve as a reliable and balanced alternative to crowd-sourced human labelers in our experiments. Additionally, we examined the eight instances where GPT-4’s labels differed from the human expert’s labels. These cases were also found to be uncertain during human labeling, suggesting that GPT-4 not only avoids individual human biases but also achieves a high level of precision comparable to human-labeled results.
Agent Prompts. We demonstrate the Agent Prompts in Table 11 for reproducibility.
Appendix B Quizzes, CoI survey, and Quality Test
@@
In this appendix section, we present detailed designs of the quizzes, surveys, and quality tests in our experiments.
Quizzes. Quizzes are used in the TAGI course as a tool to measure learning outcomes. For each chapter, the quiz assesses students’ understanding of the course concepts. The quiz questions are multiple-choice, but the number of correct answers is not disclosed to the students. To score points, students must select all the correct answers. It is important to note that the quizzes are more difficult than the quality test, and only TAGI includes quizzes, whereas both courses include the quality test. In Table 6, we present three sample questions from the quiz in TAGI’s first lecture as examples.
| 1. Which of the following statements about neural network models is correct? |
| A. Graph neural networks are designed to process image data. |
| B. The multi-head self-attention mechanism in Transformer models helps the model better capture semantic dependencies between contexts. |
| C. BERT and GPT are two classic language model architectures, with BERT being more suited for text generation tasks. |
| D. The backpropagation algorithm can only be used for shallow neural networks and is not applicable to deep neural networks. |
| 2. Why is it said that we are currently moving towards artificial general intelligence (AGI)? |
| A. Achieved architectural unification, consolidating domain-specific architectures into the Transformer architecture. |
| B. Achieved task unification, merging task-specific small models into a general large model. |
| C. Achieved modality unification, converting various modal data into character sequences. |
| D. Achieved computational efficiency unification, simplifying all computational tasks into low-cost, low-resource operations. |
| 3. Which of the following statements is correct? |
| A. Large models have already reached a performance bottleneck, and further increasing model and data size will no longer improve performance. |
| B. Training and learning in large models follow three steps: pre-training, supervised fine-tuning, and learning from human feedback, with each step requiring a large amount of manually labeled data. |
| C. AlphaGo, which defeated a human Go champion, is a classic example of artificial general intelligence. |
| D. Large models can learn to use tools like humans to complete tasks. |
@@
CoI Survey. Table 7 illustrates how the surveys in the ablation systems were structured to evaluate three crucial dimensions of the learning experience: cognitive presence, teaching presence, and social presence. Each dimension includes a detailed rating guidelines to ensure consistent and reliable feedback from diverse users.
@@
Quality tests. The quality tests, administered after participants engaged with the simulated classrooms in the ablation systems, were designed to exclude low quality data from those who didn’t participate in the course. Therefore, as shown in Table 8 and Table 9, the tests include basic concepts in the course, and are much easier than quizzes. All questions were meticulously crafted and verified by subject matter experts.
| A. Proprietary Intelligence |
| B. Symbolic Intelligence |
| C. General Intelligence |
| D. Neural Network Intelligence |
| A. Masked Language Model |
| B. Next Sentence Prediction |
| C. Possibility Memorization |
| D. Next Token Prediction |
| A. Self-supervised Pre-training |
| B. Supervised Fine-tuning |
| C. Reinforcement Learning from Human Feedback |
| D. Instruction Tuning |
| A. In-context Learning |
| B. Chain-of-Thought |
| C. Sentiment Analysis |
| D. Instruction Following |
@@
Questions are multiple-choice, some with multiple correct answers, testing whether the participants are actively engaged in the experiment.
Appendix C Supplementary Experiment Results
C.1 FIAS Matrices of Users
In addition to the FIAS matrix of the entire classroom, we also provide the FIAS matrix of human users that demonstrate their interaction patterns. As depicted in Figure 5, there is a high frequency of interaction between users, teachers, and peers. Notably, the predominant user activities involve posing questions to the teacher (interactions (5, 8) and (5, 9)) and engaging in discussions with classmates (interaction (9, 9)).
C.2 Results of Learning Outcomes of HSU
We further illustrate the results of the self-reported survey in HSU. Due to the characteristic of the course (focused on developing university-level skills), HSU employed qualitative analysis to discuss learning outcomes, in contrast to TAGI’s quantitative analysis. HSU students wrote self-learning reports after the course, and we selected several cases related to their learning outcomes, which we anonymized and presented for illustration. The cases include three main topics and capabilities in HSU: Setting Academic Development Objectives, Problem Solving, and Personal Development.
Setting Academic Development Objectives: "I am now entering my third year and am eager to begin scientific research, though I am somewhat unsure of where to start. My previous setback in a project has made me doubt my research abilities. However, after using the "Self-Assessment Scale for Innovative Potential" (Taught in HSU), I realized that, based on my performance in coursework and the project, I have already demonstrated some innovative potential. So whenever I begin to doubt my abilities, I use this to alleviate my anxiety and concerns, reminding myself that great innovations require more effort and come with greater challenges."
Problem Solving: "I switched to using an efficiency journal to track what I accomplished each day. The first day was somewhat rough, but I felt like I could see how time was flowing and what traces it left behind. Over the next four days, I began to record more meticulously, using colored pens to mark my mood, and writing a journal entry at the end of each day. I didn’t complete many tasks in a day, but my focused work time increased from 5.5 hours to 9 hours."
Personal Development: "I successfully joined my professor’s research group and became a part of the team. By utilizing scientific time management techniques such as Gantt charts, schedules, daily task lists, and the Pomodoro technique (Taught in HSU), I aim to better manage my daily time. I hope to improve my GPA a little more."
C.3 Statistical Results of Ablation Systems
We also illustrate the statistical results of each ablation systems in Table 10, including the output length of each agents and users.
| Course | Teacher | Assistant | Classmates | User |
| TAGI | 353.0 | 82.3 | 123.0 | 18.9 |
| - w/o cla. | 358.2 | 71.1 | - | 13.9 |
| - w/o int. | 398.8 | - | - | - |
| HSU | 218.3 | 90.6 | 147.7 | 15.5 |
| - w/o cla. | 212.3 | 68.2 | - | 8.2 |
| - w/o int. | 228.5 | - | - | - |
| Role | Prompt Templates |
| Teacher | [role description] You are Prof. X, a virtual AI instructor specializing in artificial intelligence courses. [behaviors] When students ask questions, you provide concise and clear answers and encourage them to continue learning. If students do not ask questions or express uncertainty, you use encouraging words to continue the lesson. For difficult questions, you suggest leaving them for later. [format] Your input is a segment of the chat history from the class; please return only the responses from your role. … |
| Assistant | [role description] As a virtual classroom teaching assistant, your main role is to provide precise supplementary information to help deepen students’ understanding of the lesson content. [behaviors] You will be very careful in choosing when to speak, ensuring that your supplements and questions are beneficial and appropriate, without repeating the teacher’s lecture or unnecessarily interrupting the course flow. … Your goal is to enhance classroom interaction and learning efficiency through concise and precise contributions while maintaining a friendly and encouraging tone. …[format] Your input is a segment of the chat history from the class … [course information] Below is information about the course, which you should use to assist your answers when users inquire about related information, ensuring the correctness of your answers:… |
| Class Clown | [role description] You are a student nicknamed ’Class Clown’ who plays the role of a student in a virtual classroom environment, interacting with teachers, students, and teaching assistants. [behaviors] You are designed to express opinions on class materials when it is your turn to speak, providing perspectives that may be humorous, insightful, or intentionally divergent, but always relevant to the topics being discussed by the teacher and students. Your goal is to enrich classroom dialogue with a blend of accuracy and fun, avoiding off-topic remarks and ensuring contributions are relevant to the course focus. You creatively engage in classroom topics, balancing knowledge and entertainment while staying on topic. [format] … |
| Deep Thinker | [role description] You are a classroom assistant named "Deep Thinker", responsible for reflecting on the current teaching content, raising counterexamples or questions to promote classroom discussion. [behaviors] Your goal is to analyze the teaching content and raise relevant and constructive counterexamples or questions. If more context or explanation is needed, feel free to ask. The counterexamples or questions should be appropriate and ensure content safety. Raise counterexamples or questions in critical thinking contexts. [format] … |
| Note Taker | [role description] The Note-Taker is a diligent student who listens to the classroom chat and extracts key information to create concise notes that summarize previous discussions and lectures. [behaviors] These notes are short, presented in a friendly, student-like tone, as if sharing with classmates. The notes emphasize quality and brevity, removing unnecessary information and focusing only on the key points, excluding course and teacher introductions. [format] … |
| Inquisitive Mind | [role description] You are a classroom student assistant named "Curious Baby", you excel at asking deep, thought-provoking questions based on the lesson content, helping students better understand and explore knowledge. [behaviors] Your questions are often unexpected, challenging, and able to spark students’ curiosity and thinking. Your chat style is lively, fun, and full of childlike wonder and curiosity, but you won’t ask questions unrelated to the lesson content. All chat content must benefit the students’ learning. [format] … |